OpenStudy (anonymous):

2. What is the lateral surface area of the prism? (1 point) 283.8 m² 292.4 m² 325.4 m² 407 m²

OpenStudy (anonymous):



OpenStudy (anonymous):

@qweqwe123123123123111

OpenStudy (anonymous):

WTH are they trying to accomplish with the term "lateral" surface area? The only answer on your list that matches anything is the TOTAL surface area of the box. And that is found by adding the area of the 2 ends, the 2 faces, and the top and the bottom surfaces. That's: 2*3*5.5 + 2*3*17.2 +2*5.5*17.2 = your answer

OpenStudy (anonymous):

That is the answer to the previous one I already did... I'm not sure about lateral either... do you know anyone who could help us?

OpenStudy (anonymous):

I just did. 2*3*5.5 + 2*3*17.2 + 2*5.5*17.2 = a number. That number is on your list of answers. I tried every other way of adding the sides up to see if some other combination matched any of your answers. They do not. Therefore, the ONLY number that's on your list that matches any kind of surface dimensions on your prism is the TOTAL surface area of your prism. And that's the formula I gave you.

OpenStudy (anonymous):

I'm sorry but that CANT be right... the previous question with the SAME diagram has that answer, they wouldn't ask the same Q twice in a row... I think the 'lateral' makes a big difference but I don't know HOW! :(

OpenStudy (anonymous):

Stand by, I'm googling stuff to see what I can find.

OpenStudy (anonymous):

Thank you SO much! :D

OpenStudy (anonymous):

This pisses me off. It turns out that the "lateral surface" is the perimeter of the base times its height. Now that had already crossed my mind in my previous attempt, and I checked that using the end faces and the front face and back face of the prism shown. IOW, that would have been (3+17.2)*2*5.5, but that only gave 192.2 which isn't on your list. I had also tried doing the surfaces over the top, around the sides and across the bottom, which was (17.2+5.5)*2*3 = 136.2, but THAT wasn't on your list either. I didn't bother doing the 3rd possibility because I thought that was absurd. And that would have been taking the "lateral surfaces" as the top, bottom, rear face, and back face of the prism in your picture. However, that's what it is. The "perimeter" of the end caps are (3+5.5)*2 = 17 The "height" now becomes it's length of 17.2, which gives us 17*17.2 = 292.4 Anyway, the secret's finally out and you have your answer. :-) Just be wary of these "lateral surfaces" on boxes, because as you can see, there's 3 different sets of them on every 6 sided prism. :-)